OpenStudy (anonymous):

Find cos θ if sin θ= -5/13 and tan θ > 0.

OpenStudy (anonymous):

choices below:

OpenStudy (anonymous):

@Hero @PFEH.1999 @SolomonZelman

hartnn (hartnn):

use the relation : \(\Large \sin^2 \theta + \cos^2 \theta =1\) you have sin, find cos

OpenStudy (anonymous):

i have to use Pythagorean theorem though

OpenStudy (anonymous):

you can use it too ;) draw a triangle and write WHAT is sin

hartnn (hartnn):

then make a right triangle, 

hartnn (hartnn):

fill a, b,c using the relation sin θ= -5/13

OpenStudy (anonymous):

Sin= opp/hyp

hartnn (hartnn):

correct